# A Novel Spike Transformer Network for Depth Estimation from Event Cameras via Cross-modality Knowledge Distillation

### v2

@@ Abstract
Depth estimation is crucial for interpreting complex environments, especially in areas such as autonomous vehicle navigation and robotics. Nonetheless, obtaining accurate depth readings from event camera data remains a formidable challenge. Event cameras operate differently from traditional digital cameras, continuously capturing data and generating asynchronous binary spikes that encode time, location, and light intensity. Yet, the unique sampling mechanisms of event cameras render standard image based algorithms inadequate for processing spike data. This necessitates the development of innovative, spike-aware algorithms tailored for event cameras, a task compounded by the irregularity, continuity, noise, and spatial and temporal characteristics inherent in spiking data.Harnessing the strong generalization capabilities of transformer neural networks for spatiotemporal data, we propose a purely spike-driven spike transformer network for depth estimation from spiking camera data. To address performance limitations with Spiking Neural Networks (SNN), we introduce a novel single-stage cross-modality knowledge transfer framework leveraging knowledge from a large vision foundational model of artificial neural networks (ANN) (DINOv2) to enhance the performance of SNNs with limited data. Our experimental results on both synthetic and real datasets show substantial improvements over existing models, with notable gains in Absolute Relative and Square Relative errors (49% and 39.77% improvements over the benchmark model Spike-T, respectively). Besides accuracy, the proposed model also demonstrates reduced power consumptions, a critical factor for practical applications.
@@ 1 Introduction
Event-based cameras are bio-inspired sensors that capture visual information asynchronously, reporting brightness changes in real-time [1, 2]. Unlike traditional cameras, the major advantages of event-based sensors include low latency between triggered events [3], low power consumption [4] and high dynamic range [5].These benefits come directly from the hardware design, and event-based cameras have being used in a variety of fields such as 3D scanning [6], robotic vision [7], and automotive industry [8]. However, in practice, the event-based sensor captures unique spike data that encodes information of light intensity changes in a scene. The noise in the data is extremely high and there is a lack of generalized processing algorithms for these data to provide comparable capabilities similar to traditional vision algorithms for data from conventional digital camera.
@@
Spiking Neural Networks (SNNs) are bio-inspired neural network models that mimic the behaviour of biological neural networks. SNNs use discrete functions called spikes to represent and process information, as opposed to continuous values used in traditional vision algorithms [9].Therefore, it is naturally an ideal paradigm for processing the output of event cameras [10, 11, 12].
@@
Depth estimation is a challenging task in computer vision, which are widely used in autonomous driving, automated robotics, agricultural growth monitoring, and forest carbon emission monitoring. Current state-of-the-art depth prediction efforts have mainly focused on combining standard frame-based cameras with artificial neural networks (ANN) [13, 14, 15]. However, event-based cameras and associated processing algorithms for depth-sensing applications are still at an infancy stage [5].A number of challenges remain, including lack of SNN backbone for feature extraction and poor SNN model performance, are two major issues.
@@
Lack of SNN backbone designed for spike data depth estimation. The event-based camera generates continuous spike streams in a binary irregular data structure that possesses ultra-high temporal features. SNN is applicable to event camera datasets and able to improve the depth estimation performance by exploiting advanced architecture of ANN, such as ResNet-like SNNs and Spiking Recurrent Neural network [16, 17, 18, 19]. Vision transformer [20, 21](ViT) which is based on self-attention mechanism to capture long-distance dependencies, especially spatial-temporal features in images/videos, is currently the most popular ANN structure. It improves the performance of AI in many computer vision tasks such as image classification/ segmentation [Strudel2021Segmenter:Arnab2021ViViT:, 22], object detection [23]and depth estimation [24, 25]. Transformer-based SNNs are a kind of new form of SNNs that combine transformer architectures with SNNs, offering great potential to break the performance bottleneck of SNNs on spike stream data. In [26], the authors used the original ViT structure as a backbone to extract features from both spatial and temporal domains in spike data. The result demonstrated the suitability of the transformer for extracting spatio-temporal features in spike data. However, the original transformer structure has a large number of multiplication operations and has excessive computational energy consumption compared to SNNs. In [27] and [28], the author proposed a pure spike driving self-attention and residual connection to avoid non-spike computations in transformer. This pushes a major step forward for the potential use of transformer for the depth estimation from spike data.
@@
SNN model performance. One of the biggest challenges with SNNs currently is their inability to achieve an equivalent performance on spiking data like ANNs do on non-spiking data. The spiking data is non-differentiable, which makes it difficult to use the backpropagation algorithm to train SNNs. Gradient-based Backpropagation is a powerful algorithm for training ANNs, but it cannot be used directly with SNNs [29]. Converting ANN to SNN directly is a solution but it may introduce errors of uncertainty or lose temporal information of spikes [19]. Meanwhile, the event-based datasets are small compared to the static images used in traditional ANN training, making SNNs prone to overfitting and limits their generalization ability [29].Knowledge distillation is a technique in deep learning by transferring knowledge from the teacher model to the student model. It allows to train a lightweight model (student model) to be as accurate as a larger model (teacher model). Currently, there are already some ANN models trained with massive data that can achieve zero-shot for depth estimation [30, 31, 32].These models logically have the potential to be transferred to the SNN model training.
@@
In this work, by leveraging the unique biological properties of SNNs and the advanced cross-modality knowledge distillation from a visual foundation model(DINOv2), we propose a spike-driven transformer for depth estimation via cross-modality knowledge distillation. To the best of our knowledge, this is the first exploration of transformer based SNN for depth estimation. We highlight three main contributions of our study as follows:
  1. 1.
    We introduce an innovative spike transformer network specifically designed for depth estimation. This network incorporates spike-driven residual learning and spike self-attention mechanisms to eliminate the need for floating-point and integer-float multiplications, adhering to the principled spike-based operation. This approach significantly reduces energy consumption while ensuring robust performance.
  2. 2.
@@
    We develop a comprehensive single-stage knowledge distillation framework, deriving insights from both the final and intermediate layers of the large vision foundation ANN model (DINOv2). Utilizing domain loss and semantic loss, our framework effectively transfers knowledge to the SNN, facilitating the efficient training on limited datasets.
  3. 3.
    We have conducted thorough experimental evaluation by comparing the proposed approach with the state-of-the-art methods on both real and synthetic datasets. The results demonstrate that our proposed method reliably predicts depth maps and outperforms competing methods by a significant margin.
@@ 2 Related works
This section covers related works in image and event based Monocular depth estimation, Spiking Neural Networks (SNNs) and Knowledge distillation for SNN.
@@ 2.1 Image-based and Event-based Monocular Depth Estimations
Depth estimation from images aims to measure the distance of each pixel relative to the camera. Monocular depth estimation, i.e. based on a single image, is a challenging but promising technology. It has the advantage of not requiring two images, which makes it more practical for applications where it is not possible to take a pair of images, such as mobile devices. Depending on the type of data used, we can divide the monocular depth estimation method into Image-based and Event-based methods [14]. The Image-based monocular depth estimation uses the information in an RGB image to estimate depth, while the event-based method uses the spike data generated by event camera. Image-based monocular depth estimation is more common than event-based monocular depth estimation because RGB images are easier to collect and process. Event cameras are a new type of sensors that output brightness changes in the form of an asynchronous "event" stream, rather than static images. This makes them well-suited for depth estimation in challenging conditions, such as low light and fast motion [Hidalgo-Carrió2020Learning].
@@
The latest developments in deep learning have made it possible to develop monocular depth estimation models that can achieved satisfactory accuracy and robustness[14, 33, 34].Similar to other deep learning models, these models typically consist of a generalized encoder that extracts abstract features from context information and a decoder that recovers depth information from the features. For RGB images, in [35], the author used ResNet-50 as an encoder and a novel up-sampling blocks as a decoder to estimate depth from a single RGB image. In [32], the authors utilised ViT instead of convolutional networks as the backbone for a depth estimation task. Experiments have found that transformer is able to provide finer and more globally consistent predictions than traditional convolutional networks. For event data, the authors [Hidalgo-Carrió2020Learning] presented a new deep learning model called E2Depth that can estimate depth from event cameras with high accuracy. A fully convolutional neural network based on the U-Net architecture [36] were used in this work. In [37], a multiscale encoder was used to extract features from mixed-density event stacking and an upscaling decoder was used to predict the depth. The transformer structure was also used in event-based monocular depth estimation. In [38], EReFormer was proposed to estimate depth from event cameras with superior accuracy based on transformer.
Traditional deep learning relies on supervised learning, which typically requires large amounts of training data. Especially for transformer architectures, which can learn long-range dependencies and have a tremendous demand for data [21]. For monocular depth estimation, with massive training data, the performance of the model can be improved[32]. The requirement for large data training significantly limits the scalability and usability of the model. Moreover, it is almost impossible to train a deep learning model in a supervised manner on the spike data stream because it is too time-dense to obtain paired depth labels.
In response to these challenges, self-supervised learning, Un-supervised learning, and knowledge distillation have emerged as new hotspots for deep learning model training on spiking data. Self/Un-supervised learning is a type of machine learning where the model learns from unlabelled data, while Knowledge distillation is a technique that transfer knowledge from large, trained model to a new model. It is reasonable to expect that these techniques have potential to be utilised in event-based monocular depth estimation. In the paper [38], the author distilled the knowledge from large trained models to improve the performance of the proposed model. [39] proposed an unsupervised spike depth estimation method via knowledge transfer from an image-based depth estimation model.
@@ 2.2 Spiking Neural Networks (SNNs)
Unlike traditional deep learning models that convey information using continuous decimal values, SNNs use discrete spike sequences to calculate and transmit information. Spiking neurons receive continuous values and convert them into spike sequences. A number of different spiking neuron models have been proposed. Hodgkin-Huxley model is one of the first models that describes the behaviour of biological neurons [40], which is a fundamental model to explain how spike flows in neurons, but the model is too complex to implement in silicon. Lzhikevich model [41]simplifie d Hodgkin-Huxley model, which is a two-dimensional model that describes the dynamics of the membrane potential of a neuron. The leaky integrate-and-fire (LIF) neuron is another simple model of a neuron that is widely used in neuroscience and artificial neural networks. It is simpler than Lzhikevich model but captures the essential features on how neurons work. It can be used to build SNN and implemented in very-large-scale integration (VLSI)[5]. The membrane potential of the LIF neuron is governed by the following equation:
$dv/dt=I-v/\operatorname{tau}$ (1)
@@
where v is the membrane potential, t is time, tau is a time constant, and I is the current. The current, I, can be either excitatory or inhibitory. Excitatory currents make the membrane potential more positive, while inhibitory currents make it more negative. When the membrane potential reaches the threshold, the neuron fires an action potential. In this work, LIF is used to build the proposed model.
In the SNN structure design, similar to ANN, as the depth of spiking neural networks (SNNs) increases, their performances have been significantly improved [16, 17, 42]. Currently, most SNNs have borrowed structures from ANNs, which can be categorized into two main groups: CNN-based SNNs and transformer-based SNNs.
@@
ResNet, as the most successful CNN model has been extensively studied to extend the depth of SNNs [16, 17]. SEW ResNet [16] overcomes the vanishing/exploding gradient problem in SNNs by using a technique called spike-timing-dependent plasticity (STDP). It has been shown effective in a variety of tasks, including image classification and object detection. However, Convolutional networks possess translation invariant and locally dependent, but their calculation has a fixed receptive field, limiting their ability to capture global dependence. In contrast, ViT is based on self-attention mechanisms that can capture long-distance dependencies. They are based on the Transformer architecture, which was originally developed for natural language processing tasks.
Transformer-based SNNs represent a novel form of SNNs that combines the transformer architecture with SNNs, providing great potential to break through the performance bottleneck of SNNs.
In Yao et al. [43], and Zhou et al[27], proposed two different Spike-Driven Self-Attention models. They utilised only mask and addition operations to avoid multiplication, which are efficient and have low computational energy consumption. Zhou et al [28], proposed Spikingformer , which modified the residual connection to purely event-driven, which is energy efficient while improving performance.
@@
Currently, ViTs have been shown to the achieve state-of-the-art results in a variety of vision tasks, including image classification, object detection, segmentation and depth estimation. However, the transformer based SNNs is still in its infancy in depth estimation[26]. There are two major issues: 1) the difficulty in training pure transformer based SNN models; 2) and the limited availability of paired depth data in event data to support transformer model training. Knowledge distillation provides a manner to train new or small models using well pretrained models. In this work, we propose a knowledge distillation method to bring ANN model knowledge into SNN.

### v3

# A Novel Energy-Efficient Spike Transformer Network for Depth Estimation from Event Cameras via Cross-modality Knowledge Distillation

@@ Abstract
Depth estimation is a critical task in computer vision, with applications in autonomous navigation, robotics, and augmented reality. Event cameras, which encode temporal changes in light intensity as asynchronous binary spikes, offer unique advantages such as low latency, high dynamic range, and energy efficiency. However, their unconventional spiking output and the scarcity of labelled datasets pose significant challenges to traditional image-based depth estimation methods. To address these challenges, we propose a novel energy-efficient Spike-Driven Transformer Network (SDT) for depth estimation, leveraging the unique properties of spiking data. The proposed SDT introduces three key innovations: (1) a purely spike-driven transformer architecture that incorporates spike-based attention and residual mechanisms, enabling precise depth estimation with minimal energy consumption; (2) a fusion depth estimation head that combines multi-stage features for fine-grained depth prediction while ensuring computational efficiency; and (3) a cross-modality knowledge distillation framework that utilises a pre-trained vision foundation model (DINOv2) to enhance the training of the spiking network despite limited data availability. Experimental evaluations on synthetic and real-world event datasets demonstrate the superiority of our approach, with substantial improvements in Absolute Relative Error (49% reduction) and Square Relative Error (39.77% reduction) compared to existing models. The SDT also achieves a 70.2% reduction in energy consumption (12.43 mJ vs. 41.77 mJ per inference) and reduces model parameters by 42.4% (20.55M vs. 35.68M), making it highly suitable for resource-constrained environments. This work represents the first exploration of transformer-based spiking neural networks for depth estimation, providing a significant step forward in energy-efficient neuromorphic computing for real-world vision applications.
keywords:
Spiking Neural Network (SNN), Event Camera, Transformer, Depth Estimation, Knowledge Distillation, Neuromorphic Computing, Energy-Efficient Computing.
journal: Neural Networks
@@ 1 Introduction
Event-based cameras are bio-inspired sensors that capture visual information asynchronously, reporting brightness changes in real-time (Gallego et al., 2022; Tayarani-Najaran and Schmuker, 2021). Due to hardware design, they offer many advantages over traditional cameras, including low latency between trigger events (Chen and Guo, 2019), low power consumption (Posch et al., 2011), and high dynamic range (Furmonas et al., 2022). These event-based sensors are already being used in a variety of fields such as 3D scanning (Huang et al., 2022), robotic vision (Cao et al., 2021), and the automotive industry(Chen et al., 2020). In practice, event-based sensors capture unique spiking data, encoding information about changes in light intensity in a scene. The noise in the data is extremely high and there is a lack of generalized processing algorithms to provide capabilities comparable to traditional vision algorithms for conventional digital camera data.
@@
Spiking Neural Networks (SNNs) are bioinspired neural network models that mimic the behavior of biological neural networks. SNNs use discrete functions called spikes to represent and process information, as opposed to continuous values used in traditional vision algorithms (Rafi, 2021). Therefore, it is naturally an ideal paradigm for processing the output of event cameras (Lee et al., 2020; Auge et al., 2021; Cordone et al., 2021).
@@
Depth estimation is a challenging task in computer vision, used widely in autonomous driving, automated robotics, agricultural growth monitoring, and forest carbon emission monitoring. Current state-of-the-art depth prediction efforts have mainly focused on combining standard frame-based cameras with artificial neural networks (ANN) (Zhao et al., 2020; Ming et al., 2021; Laga et al., 2022). Event-based cameras and associated processing algorithms for depth-sensing applications are still in their infancy (Furmonas et al., 2022). Two major challenges include the lack of a SNN backbone for feature extraction and poor SNN model performance.
@@
The lack of SNN backbone designed for spike data depth estimation. The event-based camera generates continuous spike streams in a binary irregular data structure that possesses ultrahigh temporal features. SNNs are applicable to event camera datasets and able to improve the depth estimation performance by exploiting advanced architecture of ANN, such as ResNet like SNNs and Spiking Recurrent Neural network (Fang et al., 2021; Hu et al., 2023; Yin et al., 2020; Diehl et al., 2016). Vision transformer (Dosovitskiy et al., 2020; Han et al., 2023)(ViT), is currently the most popular ANN structure and is based on a self-attention mechanism to capture long-distance dependencies, especially spatial-temporal features in images/videos. It improves the performance of AI in many computer vision tasks such as image classification/ segmentation (Strudel et al., 2021; Arnab et al., 2021; Liu et al., 2021), object detection (Sun et al., 2021) and depth estimation (Yang et al., 2022; Zhao et al., 2022). Transformer-based SNNs are a new form of SNN combining transformer architectures with spiking neurons, offering great potential to break the performance bottleneck on spike stream data. In Zhang et al. (2022), the authors used the original ViT structure as a backbone to extract features from both spatial and temporal domains in spike data. The result demonstrated the suitability of the transformer for extracting spatio-temporal features. However, the original transformer structure has a large number of multiplication operations and excessive computational energy consumption compared to SNNs. In Zhou et al. (2022) and Zhou et al. (2023), the author proposed a pure spike driven self-attention and residual connection to avoid non-spike computations. This was a major step forward in the potential use of transformer for depth estimation from spike data.
@@
SNN model performance. One of the biggest challenges with SNNs currently is their inability to achieve equivalent performance on spiking data to ANNs on non-spiking data. Gradient-based backpropagation is a powerful algorithm for training ANNs, but as spiking data is non-differentiable it cannot be used directly with SNNs (He et al., 2023). Converting ANN to SNN is a solution but it may introduce errors of uncertainty or lose the temporal information of spikes (Diehl et al., 2016). Meanwhile, the number of event-based datasets are small compared to the static images used in traditional ANN training, making SNNs prone to overfitting and limiting their generalization ability (He et al., 2023). Knowledge distillation is a technique in deep learning to transfer knowledge from the teacher model to the student model. It allows training of a lightweight model (student model) to be as accurate as a larger model (teacher model). Currently, there are already some ANN models trained with massive data that can achieve zero-shot for depth estimation (Oquab et al., 2023; Ranftl et al., 2020, 2021). Logically, the accuracy of these models has the potential to be transferred to the SNN model during training.
@@
In this work, we propose a novel energy-efficient spike transformer network for depth estimation, leveraging cross-modality knowledge distillation to combine the biological efficiency of SNNs with the advanced feature extraction capabilities of a visual foundation model (DINOv2). To the best of our knowledge, this is the first exploration of a transformer-based SNN for depth estimation, marking a significant advancement in the field. The proposed framework comprises three key components, each contributing uniquely to its overall effectiveness:
  1. 1)
    We introduce a novel energy-efficient spike-driven transformer that eliminates conventional floating-point operations through carefully designed spike-based attention and residual mechanisms. This network incorporates two essential components: a spiking patch embedding module that converts raw event data into spike-based tokens while preserving temporal-spatial information, and spiking transformer blocks that integrate Spiking Self-Attention (SSA) and Spiking MLP for efficient feature processing. This design significantly reduces energy consumption while ensuring robust performance.
  2. 2)
    We develop a fusion depth estimation head that combines features from multiple transformer stages to enable fine-grained depth prediction while maintaining efficient computation. Unlike conventional CNN-based methods that rely on downsampling, our fusion head preserves critical spatial information through a multi-scale feature integration approach, proving particularly effective in challenging scenarios such as low-light conditions and distant object detection.
  3. 3)
@@
    We propose a single-stage cross-modality knowledge distillation framework that leverages a large vision foundation model (DINOv2) to enhance SNN training with limited data. By utilis ing domain loss and semantic loss, our framework effectively transfers knowledge from both final and intermediate layers of DINOv2 to the spike-driven transformer.
@@ 2 Related works
This section presents a literature review of existing research in monocular depth estimation, SNNs, and knowledge distillation, highlighting the key challenges that motivate our work.
@@ 2.1 Image-based and Event-based Monocular Depth Estimations
Depth estimation from images aims to measure the distance of each pixel relative to the camera. Monocular depth estimation is a challenging but promising technology. It has the advantage of only requiring one image unlike traditional depth estimation, which makes it more practical for applications where it is not possible to take a pair of images, such as mobile devices. Depending on the type of data used, we can divide monocular depth estimation into Image-based and Event-based methods (Ming et al., 2021). Image-based monocular depth estimation is more common as it estimates depth using the information in RGB images, which are easy to collect and process. The event-based method uses spike data generated by an event camera, a new type of sensor that outputs brightness changes in the form of an asynchronous ”event” stream, rather than static images. This makes them well-suited for depth estimation in challenging conditions, such as low light and fast motion (Cordone et al., 2021), but the data is harder to collect and process.
@@
The latest developments in deep learning have made it possible to develop monocular depth estimation models that can achieve satisfactory accuracy and robustness (Ming et al., 2021; Khan et al., 2020; Li et al., 2020). Similar to other deep learning models, these models typically consist of a generalized encoder that extracts abstract features from context information and a decoder that recovers depth information from the features. For RGB images, in Laina et al. (2016), the author used ResNet-50 as an encoder and novel up-sampling blocks as a decoder to estimate depth from a single RGB image. In Laina et al. (2016), the authors utilised ViT instead of convolutional networks as the backbone for a depth estimation task. Experiments have found that transformer is able to provide finer and more globally consistent predictions than traditional convolutional networks. For event data, the research is still in its infancy. The authors (Cordone et al., 2021) presented a new deep learning model called E2Depth that can estimate depth from event cameras with high accuracy. A fully convolutional neural network based on the U-Net architecture (Ronneberger et al., 2015) was used in this work. In Nam et al. (2022), a multiscale encoder was used to extract features from mixed-density event stacking and an upscaling decoder was used to predict the depth. The transformer structure has also been used in event-based monocular depth estimation. In Liu et al. (2022b), EReFormer was proposed to estimate depth from event cameras with superior accuracy based on transformer.
However, these models predominantly utilise traditional deep learning frameworks, overlooking the unique potential of event-based data. Existing research identifies two key challenges that remain unaddressed. The first challenge lies in the unique characteristics of event camera data. These cameras produce continuous streams of sparse, asynchronous data, which traditional frame-based methods struggle to process efficiently. They also provide very high temporal resolution, capturing subtle movements and rapid changes in brightness. This requires algorithms that can process data in real-time and maintain temporal accuracy (Gallego et al., 2022; Tayarani-Najaran and Schmuker, 2021). SNNs are biologically inspired networks that process information using spikes, similar to the human brain, and are well-suited to handle the event-driven nature of these cameras (Tavanaei et al., 2019). In this paper, we review the current SNN methods in the literature and propose a novel spike transformer network for depth estimation from event camera data. The second challenge is the scarcity of spike training data. High-quality, labelled datasets tailored for SNNs, particularly for tasks like depth estimation, remain limited. The acquisition and labelling of event-based data are both complex and resource-intensive, further constraining the availability of training resources. To address this limitation, knowledge distillation offers a promising solution. This involves transferring knowledge from a well-trained artificial neural network (ANN). The ANN acts as a ”teacher”, guiding the SNN, or ”student”, to learn effectively with limited event-based data.
@@ 2.2 Spiking Neural Networks (SNNs)
Unlike traditional deep learning models that convey information using continuous decimal values, SNNs use discrete spike sequences to calculate and transmit information. Spiking neurons receive continuous values and convert them into spike sequences. A number of different spiking neuron models have been proposed. The Hodgkin-Huxley model is one of the first models that describes the behaviour of biological neurons (Hodgkin and Huxley, 1952), and is fundamental to explain how spikes flow in neurons, but the model is too complex to implement in silicon. The Lzhikevich model (Izhikevich, 2003), which simplifies the Hodgkin-Huxley model, is a two-dimensional model that describes the dynamics of the membrane potential of a neuron. The leaky integrate-and-fire (LIF) neuron is another simple neuron model that is widely used in neuroscience and SNNs. It is simpler than Lzhikevich model but captures the essential features on how neurons work. It can be used to build SNNs and implemented in very-large-scale integrations (VLSI)(Hsieh and Tang, 2012). The membrane potential of the LIF neuron is governed by the following equation:
$dv/dt=I-v/\tau$ (1)
@@
where v is the membrane potential, t is time, $\tau$ is a time constant, and I is the current. The current, I, can be either excitatory or inhibitory. Excitatory currents make the membrane potential more positive, while inhibitory currents make it more negative. When the membrane potential reaches the threshold, the neuron fires an action potential. In this work, LIF is used to build the proposed model.
@@
Similar to ANNs, as the depth of SNNs increases, their performance significantly improves (Fang et al., 2021; Hu et al., 2023; Zheng et al., 2020). Currently, most SNNs have borrowed structures from ANNs, which can be categorised into two main groups: CNN-based SNNs and transformer-based SNNs. ResNet, as the most successful CNN model has been extensively studied to extend the depth of SNNs (Fang et al., 2021; Hu et al., 2023). SEW ResNet (Fang et al., 2021) overcomes the vanishing/exploding gradient problem in SNNs by using a technique called spike-timing dependent plasticity (STDP). It has been shown effective in a variety of tasks, including image classification and object detection. However, Convolutional networks possess translation invariant and locally dependent, but their calculation has a fixed receptive field, limiting their ability to capture global dependence. In contrast, ViT is based on self-attention mechanisms that can capture long-distance dependencies. They are based on the Transformer architecture, which was originally developed for natural language processing tasks. Transformer-based SNNs represent a novel form of SNNs that combines the transformer architecture with SNNs, providing great potential to break through the performance bottleneck of SNNs. Yao et al. (2023), and Zhou et al. (2022), proposed two different Spike-Driven Self-Attention models. To avoid multiplication, they utilised only mask and addition operations, which are efficient and have low computational energy consumption. Zhou et al. (2023), proposed Spikingformer, modifying the residual connection to be purely event-driven, making it energy efficient while improving performance.
@@
Currently, ViTs have been shown to the achieve state-of-the-art results in a variety of vision tasks, including image classification, object detection, segmentation, and depth estimation. However, transformer based SNNs are still in their infancy regarding depth estimation (Zhang et al., 2022). There are two major issues: 1) the difficulty in training pure transformer based SNN models; 2) the limited availability of paired depth data in event data to support transformer model training. Knowledge distillation provides a manner to train new or small models using well pretrained models. In this work, we propose a knowledge distillation method to bring ANN model knowledge into SNN.